OpenStudy (anonymous):

Help me simplify: (1/a) + (1/b) / (1/a^2) - (1/b^2) [(2^-1 + 2(3/4)^-2)/ (8^-1 - 4^-2)] * 33/37^-1

OpenStudy (anonymous):

\[\frac{ \frac{1}{a} -\frac{1}{b}}{ \frac{1}{a^2} -\frac{1}{b^2} }\] = ?

OpenStudy (anonymous):

And yes, there are negative exponents.

Directrix (directrix):

Before we start, do you know the pattern for factoring the difference of two squares. I'm talking about something like x^2 - y^2 = (x + y)*(x - y).

OpenStudy (anonymous):

Yes I do ^^ my teacher said to find the common denominator which is a^2b^2 for the denominator.

OpenStudy (anonymous):

and for the numerator it is ab.

Directrix (directrix):